# KG-QAGen: A Knowledge-Graph-Based Framework for Systematic Question Generation and Long-Context LLM Evaluation

### v1

@@ Abstract
The increasing context length of modern language models has created a need for evaluating their ability to retrieve and process information across extensive documents. While existing benchmarks test long-context capabilities, they often lack a structured way to systematically vary question complexity. We introduce KG-QAGen (Knowledge-Graph-based Question-Answer Generation), a framework that (1) extracts QA pairs at multiple complexity levels (2) by leveraging structured representations of financial agreements (3) along three key dimensions – multi-hop retrieval, set operations, and answer plurality – enabling fine-grained assessment of model performance across controlled difficulty levels. Using this framework, we construct a dataset of 20,139 QA pairs (the largest number among the long-context benchmarks) and open-source a part of it. We evaluate 13 proprietary and open-source LLMs and observe that even the best-performing models are struggling with set-based comparisons and multi-hop logical inference. Our analysis reveals systematic failure modes tied to semantic misinterpretation and inability to handle implicit relations.
@@ 1 Introduction
The increasing context length of recent language models has led to growing interest in evaluating their capabilities. Researchers have studied challenges like the Lost in the Middle problem (Liu et al., 2024b), examined the limiting factors of scaling models to long context (such as the constraints imposed by RoPE’s base value (Men et al., 2024)), and developed techniques to extend context length efficiently (Hooper et al., 2024; Ding et al., 2024). Others have focused on improving inference efficiency (Jiang et al., 2024b; Tang et al., 2024) or enhancing long-context utilization and mitigating information loss (Zhang et al., 2024d; Lin et al., 2024; Zhang et al., 2024c).
Numerous benchmarks have been developed to evaluate long-context LLMs, covering a variety of tasks and capabilities. Some benchmarks support multi-lingual evaluation (Agrawal et al., 2024; Bai et al., 2024), include tasks such as segment sorting (Wang et al., 2024a; Dong et al., 2024), numerical reasoning (Zhao et al., 2024c; Reddy et al., 2024), timeline reordering (Zhang et al., 2024a). However, certain aspects of understanding long-context remain underexplored in existing benchmarks.
Figure 1: Overview of KG-QAGen. Credit agreements are annotated to identify entities and their relationships, forming a knowledge graph representation. This graph is then used to systematically extract multi-level QA pairs, which serve as the basis for benchmarking long-context LLMs.
Real-World Documents
Understanding large, real-world documents presents a unique challenge due to complex entity relationships, co-reference variations, and cross-referenced information scattered across sections (Zaporojets et al., 2022). This complexity makes credit agreements a precise fit for long-context evaluation, as they contain networks of entities referenced in varied ways across clauses (Section 3.1) and introduce numerous challenging edge cases (Appendix A).
Question Complexity Levels
Long-context benchmarks can offer deeper insight by distinguishing question difficulty levels, enabling analysis of how performance degrades with increasing logical inference depth – whether due to retrieval limits, context length, or multi-step inference. While some benchmarks contain tasks of varying difficulty (Gupta et al., 2024; Zhao et al., 2024c; Zhu et al., 2024), none explicitly categorize question complexity levels. Moreover, our benchmark not only defines these levels but also provides a systematic approach to creating granular questions, leveraging our annotations to deterministically construct QA pairs with clearly defined complexity directions.
We introduce KG-QAGen (Knowledge-Graph-based Question-Answer Generation), a framework for structured extraction of QA pairs from manually annotated documents via knowledge-graph representations. Our approach structures questions along three dimensions: the number of hops, the use of set operations, and whether the answer is singular or plural. Using a knowledge-graph-inspired representation, we create diverse questions and extract answers efficiently with SPARQL queries. This structured approach ensures scalability, flexibility, and controlled dataset creation, enabling a more precise evaluation of long-context LLMs.
We apply KG-QAGen to 170 credit agreements, resulting in KG-QAGen-D, a dataset of 20,139 QA pairs categorized by three complexity levels. We use it to evaluate 13 long-context LLMs and find that even on simple factual queries, some models fail to extract basic information; performance further drops with multi-hop paths and compositional questions involving set operations (Section 4.3). Our analysis shows that set-based comparisons are especially challenging, with high rates of “Not Found" responses and semantic misinterpretation (Section 5.2).
Our contributions are as follows: (1) We present a framework grounded in knowledge graphs, enabling systematic extraction of QA pairs at varying complexity levels. (2) We open-source a portion of the KG-QAGen-D to encourage future work (see Appendix F for details). (3) We release our inference pipeline and benchmarking codes to facilitate reproducible evaluation of long-context models.
2 Related Work
Table 1: Comparison of KG-QAGen-D with existing long-context LLM benchmarks. Our benchmark stands out by evaluating question complexity on a granular level, incorporating real-world SEC credit agreements, and supporting multi-hop and context-aware retrieval.
Real-world Multi-hop Context-aware Ques. complexity
Dataset documents retrieval retrieval levels Document Type # Samples
mLongRR (Agrawal et al., 2024) partially BBC News unspecified
L-Eval (An et al., 2024) partially Earnings Calls + Misc 2,008
@@ 2 Related Work
Marathon (Zhang et al., 2024a) Prior Datasets 1,530
@@
M4LE (Kwan et al., 2024) partially News + Misc unspecified
LooGLE (Li et al., 2024) partially arXiv + Misc 6,448
BAMBOO (Dong et al., 2024) partially ACL Anthology + Misc 1,502
LongBench (Bai et al., 2024) Previous Datasets 4,750
MMLongBench-DOC (Ma et al., 2024) partially arXiv + Misc 1,082
Loong (Wang et al., 2024b) SEC + Misc 1,600
FinDVer (Zhao et al., 2024b) SEC 10-K & 10-Q 2,100
DocFinQA (Reddy et al., 2024) SEC 10-K & 10-Q 7,437
@@
DocMath-Eval (Zhao et al., 2024c) partially SEC 10-K & 10-Q 4,000
KG-QAGen-D (ours) SEC Credit Agreements 20,139
Structured retrieval from long-context documents remains a key challenge, especially when handling complex entities and their relationships (Lin et al., 2025; Guo et al., 2024; Saad-Falcon et al., 2023). Multi-hop retrieval and context-aware models are particularly relevant for real-world documents.
Multi-Hop Retrieval
Multi-hop retrieval involves retrieving multiple interconnected pieces of evidence across different sections of a document or multiple documents to answer a question. This capability is crucial for handling complex queries requiring multi-step retrieval, such as tracing relationships between entities in legal documents or extracting financial conditions scattered across an agreement (Wang et al., 2025; Sarmah et al., 2024). Some existing benchmarks (An et al., 2024; Wang et al., 2024a) focus on long-context understanding without incorporating explicit multi-hop retrieval challenges.
Context-Aware Retrieval
Context-aware retrieval refers to tasks where models retrieve relevant information to answer questions based on document content rather than artificial information placement. Some existing benchmarks focus on the Needle in a Haystack problem (Agrawal et al., 2024), where small pieces of information are inserted into long texts and then retrieved. This type of retrieval does not reflect real-world scenarios where understanding the document’s structure and semantics is crucial (Munikoti et al., 2023; Lai et al., 2024).
3 The KG-QAGen Benchmark
KG-QAGen systematically extracts structured question-answer pairs across varying complexity levels for long-context evaluation. Using credit agreements (Section 3.1) annotated with a structured label schema (Section 3.2, Appendix A), we construct a knowledge graph (Section 3.3) to enable systematic query construction. Questions are designed along three dimensions – multi-hop retrieval, set operations, and answer plurality (Section 3.4) – ensuring scalable, diverse, and automated question extraction (Section 3.5). The dataset balances complexity and document coverage, making it a robust benchmark for assessing long-context retrieval.
@@ 3.1 Credit Agreements
We sampled 17 credit agreements per year from 2013 to 2022 from SEC EDGAR, totaling 170 documents. Table 2 summarizes their length by token and sentence count. Their suitability for long-context LLM benchmarking is discussed below, with supporting examples in Appendix A.
Rich in Entities and Relations
Credit agreements contain multiple organizations, individuals, roles, and locations, with entities appearing in varied forms throughout the document. Some roles have sub-roles (e.g., a lender can be a Swing-Line Lender or Lead Lender). Extracting information requires cross-referencing sections, making retrieval a challenging task.
Ambiguous Co-reference
Entities are often referenced in multiple ways; for example, “Bank of America, N.A.” may also appear as “Bank of America”. Extracting the exact full name can be challenging, as abbreviations or partial mentions may appear throughout the document.
Structured but not Standardized
While they follow a general structure (e.g., loan terms, covenants, defaults), variations in phrasing and formatting mean that the same type of information can appear in different sections or be expressed in diverse ways, preventing models from relying on fixed templates and requiring deeper understanding of the document’s content.
Real-World Significance
Credit agreements outline multi-million-dollar transactions and define the rights and responsibilities of involved parties. Accurate analysis of these documents is essential for financial institutions, legal teams, and regulators to assess risk, ensure compliance, and avoid costly misunderstandings. This makes them a highly relevant and impactful target for developing and evaluating QA systems.
@@ 3.2 Annotations
Table 2: Statistics of the source documents and their structured annotations used to construct KG-QAGen-D. The table summarizes token and sentence counts, as well as the distribution of annotated entities and relations per document. The low minimum values are due to an edge case discussed in Appendix A.3.
@@
Person Name 2 12.55 88
Person Position 2 13.20 88
Location 1 5.62 168
Location Type 1 0.43 7
@@
Org Name $\rightarrow$ Person Name 2 0.97 14
Person Name $\rightarrow$ Person Position 2 3.04 54
Org Name $\rightarrow$ Location 1 5.64 178
Location $\rightarrow$ Location Type 1 0.45 18
Our annotation schema captures key entities and relationships in credit agreements, defining seven label types covering persons, organizations, roles, and locations to represent their structure. Appendix A outlines label definitions, placement rules, and relationship guidelines, while Table 2 illustrates entity relationships and their frequency across documents.
We annotated the extracted credit agreements using Label-Studio [^1] https://github.com/HumanSignal/label-studio, supporting entity linking and relation extraction for .html files. The annotation was carried out by annotators with backgrounds in Computer Science or Mathematics, under the direct supervision of a Master of Finance and a PhD student advised by a Chair Professor of Finance, also a co-author. The annotation followed a structured, iterative approach to ensure consistency and resolve edge cases. The team first established label definitions and completed an initial annotation round. The results were reviewed to identify and discuss edge cases, which led to refinements in the annotation guidelines. To enhance reliability, two additional rounds of annotation cleaning were conducted, with annotators reviewing documents they had not previously worked on.
3.3 Knowledge Graph Representation
@@
To enable systematic and scalable QA extraction, we transform annotated credit agreements into a knowledge graph representation. This graph encodes entities of people, organizations, roles, and locations, along with their relationships, in a structured and queryable form. Figure 2 illustrates a graph which captures that Jane Smith holds the position of Vice President at ABCDE Inc. – an organization that serves as a Lender as well as an Administrative and Documentation Agent, and has its Headquarters at 10 Peachtree.... By abstracting away document-specific phrasing while preserving these logical links, the graph supports consistent interpretation across agreements and forms the basis for creating questions of varying complexity through structured queries.
We construct the knowledge graph using the Resource Description Framework [^2] https://www.w3.org/TR/rdf11-concepts/ (RDF), which represents data as subject-predicate-object triples, forming a directed labeled graph for knowledge representation. Stored in Turtle (*.ttl) format for readability, this structure ensures scalability, and efficient querying across documents using SPARQL [^3] https://www.w3.org/TR/sparql11-query/. See Appendix B for knowledge graph construction details.
@@ 3.4 Question Templates and Complexity Dimensions
Number of Hops
Set Operations
Plurality
Moving from expecting a singular answer to allowing multiple valid answers, e.g., “What companies are the [(Org Sub-Role +) Org Role] in the agreement?" if multiple roles are held.
Table 3: This table defines the question templates used to construct KG-QAGen-D, structured along three dimensions: plurality (P), number of hops (H), and set operations (#SO). It includes example templates, corresponding knowledge graph query paths, and logical operations involved. These dimensions are used to compute the overall complexity level for each QA pair.
P H #SO Example Template Hop Path Set Operation
@@ 3.5 Extraction of QA Pairs from Templates
Table 3 illustrates how questions are systematically constructed by incrementally increasing complexity across three dimensions introduced in Section 3.4: number of hops (H), plurality (P), and number of set operations (#SO). Each row in the table corresponds to a specific question template, with associated values for P, H, and #SO, the query path in the knowledge graph, and the set operation logic used.
Question instances are constructed by dynamically substituting entities into these templates using SPARQL queries over the knowledge graph described in Section 3.3. This ensures precise and scalable QA pair creation, with answer labels grounded in annotated relationships. While our framework supports arbitrarily complex combinations, we observed that highly compositional templates often become inapplicable to real documents. For instance, there are no individuals in an agreement who hold three distinct positions or no organization with the combination of roles required to fulfill a complex query. Thus, we limit template depth since such complex conditions are typically not present in the source documents. All the created templates can be found in Appendix C.
To support evaluation, we define a composite question complexity level $L\!=\!H\!+\!P\!+\!\#SO$, and group QA pairs into three categories. Easy ($L\!=\!1$): simple, single-hop, singular-answer queries. Medium ($42\!\leqslant\!L\!\leqslant\!$ 4): questions with moderate complexity via hops, set operations, or plurality. Hard ($L\!=\!5$): questions with the most complex combinations of all three dimensions.
The resulting dataset, KG-QAGen-D, comprises 20,139 QA pairs from 170 credit agreements, surpassing existing long-context QA benchmarks in size (Table 1). This scale is achieved through efficient question extraction from structured annotations using reusable templates.
@@ 4 Evaluation
In this section, we describe the setup for our experiments. We evaluate multiple long-context LLMs (listed in Section 4.1) across different question complexity levels, as defined in Section 3.4. Performance is measured using four metrics detailed in Section D.2. In addition, we conduct human evaluation of LLM responses on a small sample of QA pairs (see Section D.4). Our prompting strategy is outlined in Section D.1. The results of our level-based evaluation are presented in Table 4, with the analysis provided in Section 5.

### v2

# KG-MuLQA: A Framework for KG-based Multi-Level QA Extraction and Long-Context LLM Evaluation

@@ Abstract
We introduce KG-MuLQA (Knowledge-Graph-based Multi-Level Question-Answer Extraction): a framework that (1) extracts QA pairs at multiple complexity levels (2) along three key dimensions – multi-hop retrieval, set operations, and answer plurality, (3) by leveraging knowledge-graph-based document representations. This approach enables fine-grained assessment of model performance across controlled difficulty levels. Using this framework, we construct a dataset of 20,139 QA pairs based on financial credit agreements and evaluate 16 proprietary and open-weight Large Language Models, observing that even the best-performing models struggle with set-based comparisons and multi-hop reasoning over long contexts. Our analysis reveals systematic failure modes tied to semantic misinterpretation and inability to handle implicit relations.
KG-MuLQA: A Framework for KG-based Multi-Level QA Extraction
and Long-Context LLM Evaluation
Nikita Tatarinov,🖂 Vidhyakshaya Kannan,* Haricharana Srinivasa,* Arnav Raj, Harpreet Singh Anand, Varun Singh, Aditya Luthra, Ravij Lade, Agam Shah,🖂 Sudheer Chava Georgia Institute of Technology 🖂 Corresponding Authors: {ntatarinov3, ashah482}@gatech.edu * Indicates equal contribution
@@ 1 Introduction
Figure 1: Overview of KG-MuLQA. Credit agreements are annotated to identify entities and their relationships, forming a knowledge graph representation. This graph is then used to systematically extract multi-level QA pairs, which serve as the basis for benchmarking long-context LLMs.
Table 1: Comparison of KG-MuLQA-D with existing long-context LLM benchmarks. Multi-hop Reasoning marks whether the dataset requires reasoning across multiple evidence pieces; KG-MuLQA-D explicitly defines and scales multi-hop paths through its knowledge-graph structure, enabling controlled evaluation of relational inference depth. Systematic Q-Extraction denotes whether questions are generated deterministically rather than manually or by LLMs; KG-MuLQA-D automates QA generation from annotated graphs, preserving real document semantics without synthetic injection. Question Complexity shows whether the benchmark categorizes questions by difficulty; KG-MuLQA-D introduces a principled, multi-dimensional notion of complexity (hops + set operations + plurality), allowing granular performance analysis across reasoning levels.
Long Multi-hop Systematic Question
Dataset context reasoning Q-extraction complexity Document Type # Samples
L-Eval (An et al., 2024) Earnings Calls + Misc 2,008
@@
Marathon (Zhang et al., 2024a) partially Prior Datasets 1,530
MMLongBench-DOC (Ma et al., 2024) partially arXiv + Misc 1,082
DocFinQA (Reddy et al., 2024) partially SEC 10-K & 10-Q 7,437
@@
LooGLE (Li et al., 2024) arXiv + Misc 6,448
BAMBOO (Dong et al., 2024) ACL Anthology + Misc 1,502
LongBench (Bai et al., 2024) Previous Datasets 4,750
Loong (Wang et al., 2024) SEC + Misc 1,600
FinDVer (Zhao et al., 2024b) SEC 10-K & 10-Q 2,100
M4LE (Kwan et al., 2024) partially Prior Datasets unspecified
Michelangelo (Vodrahalli et al., 2024) partially partially synthetic unspecified
@@
mLongRR (Agrawal et al., 2024) needles BBC News unspecified
@@
RULER (Hsieh et al., 2024) needles partially SEC 10-K & 10-Q 30,000
BABILong (Kuratov et al., 2024) partially partially Books + Misc 60,000
MuSiQue (Trivedi et al., 2022) Wikipedia 24,814
KG-MuLQA-D (ours) SEC Credit Agreements 20,139
The increasing context length of recent Large Language Models (LLMs) has led to growing interest in evaluating their capabilities. Researchers have studied challenges like the Lost in the Middle problem (Liu et al., 2024b), examined the limiting factors of scaling models to long context such as the constraints imposed by RoPE’s base value (Men et al., 2024), and developed techniques to extend context length efficiently (Hooper et al., 2024; Ding et al., 2024). Others have focused on improving inference efficiency (Jiang et al., 2024b; Tang et al., 2024) or enhancing long-context utilization and mitigating information loss (Zhang et al., 2024d; Lin et al., 2024; Zhang et al., 2024c). Table 1 highlights numerous benchmarks developed to evaluate long-context LLMs on different tasks.
We introduce the Knowledge-Graph-based Multi-Level Question-Answer Extraction (KG-MuLQA) framework, outlined in Figure 1. Our approach uses knowledge-graph document representation to programmatically extract question-answer pairs of varying complexity from gold annotations. We define question templates and determine their complexity through three dimensions: the number of hops, the use of set operations, and the plurality of the answer. This structured approach ensures scalability, flexibility, and controlled dataset construction, enabling a more precise evaluation of long-context LLMs.
We apply KG-MuLQA to 170 credit agreements, resulting in KG-MuLQA-D, a dataset of 20,139 QA pairs categorized by five complexity levels. We use it to evaluate 16 long-context LLMs and find that even on simple factual queries, some models fail to extract basic information; performance further drops with multi-hop paths and questions involving set operations. Our analysis shows that set-based comparisons are especially challenging, with high rates of “Not Found” responses and semantic misinterpretation.
Our contributions are as follows. (1) We present a framework based on knowledge graphs, enabling systematic extraction of QA pairs at varying complexity levels. (2) We release a portion of the constructed dataset [^1] https://huggingface.co/datasets/gtfintechlab/KG-MuLQA-D to encourage future work (see Appendix G for details). (3) We release our inference pipeline and benchmarking codes [^2] https://github.com/gtfintechlab/KG-MuLQA to facilitate reproducible evaluation of long-context models.
2 The KG-MuLQA Benchmark
KG-MuLQA systematically extracts structured QA pairs across varying complexity levels for long-context evaluation. Using credit agreements (Section 2.1) annotated with a structured label schema (Section 2.2, Appendix A), we construct document-level knowledge graphs (Section 2.3) to enable systematic query construction. Questions are designed along three dimensions – multi-hop reasoning, set operations, and answer plurality (Section 2.4) – ensuring scalable, diverse, and automated question extraction (Section 2.5). The dataset balances complexity and document coverage (Appendix A.6), making it a robust long-context LLM benchmark.
@@ 2.1 Credit Agreements
We sampled 17 credit agreements per year from 2013 to 2022 from SEC EDGAR, totaling 170 documents. Table 2 includes their length by token and sentence count. Their suitability for long-context LLM benchmarking is detailed in Appendix A.1.
@@ 2.2 Annotations
Our annotation schema captures key entities and relationships in credit agreements, defining seven label types covering persons, organizations, roles, and locations to represent their structure. We outline label definitions in Appendix A.2, placement rules and relationship guidelines in Appendix A.3, while Table 2 includes the frequency statistics for each entity and relation type across documents. Each document was independently annotated by 3 annotators. Between the second and third (final) rounds of annotating, the inter-annotator agreement reached 80.76%, indicating substantial consistency (see Appendix A.5 for annotation process details).
2.3 Knowledge Graph Representation
To enable systematic and scalable QA extraction, we transform annotated credit agreements into knowledge graph representations. These graph encode entities of people, organizations, roles, and locations, along with their relationships, in a structured and queryable form. Figure 2 illustrates a graph which captures that Jane Smith holds the position of Vice President at ABCDE Inc. – an organization that serves as a Lender as well as an Administrative and Documentation Agent, and has its Headquarters at 10 Peachtree.... By abstracting away document-specific phrasing while preserving these logical links, the graphs support consistent interpretation across agreements and form the basis for creating questions of varying complexity through structured queries.
Table 2: Statistics of the source documents and their structured annotations used to construct KG-MuLQA-D. The table summarizes token and sentence counts, as well as the distribution of annotated entities and relations per document. The low minimum values are due to an edge case discussed in Appendix A.4.
@@ 2.3 Knowledge Graph Representation
Person Name 0 12.55 88
Person Position 0 13.17 88
Location 0 5.61 168
Location Type 0 0.45 7
@@
Org Name $\rightarrow$ Person Name 2 12.55 88
Org Name $\rightarrow$ Person Position 0 13.17 88
Person Name $\rightarrow$ Person Position 0 12.93 86
Org Name $\rightarrow$ Location 0 6.32 178
Location $\rightarrow$ Location Type 0 0.45 18
@@
We construct knowledge graphs using the Resource Description Framework [^3] https://www.w3.org/TR/rdf11-concepts/ (RDF), which represents data as subject-predicate-object triples, forming a directed labeled graph for knowledge representation. Stored in Turtle (*.ttl) format for readability, this structure ensures scalability and efficient querying across documents using SPARQL [^4] https://www.w3.org/TR/sparql11-query/ (see Appendix B for KG construction details).
@@ 2.4 Question Templates and Complexity Dimensions
Table 3: This table illustrates the question templates used to construct KG-MuLQA-D, structured along three dimensions: plurality (P), number of hops (H), and set operations (#SO). It includes example templates, corresponding knowledge graph query paths, and logical operations involved. These dimensions are used to compute the overall complexity level for each QA pair. The full list of templates can be found in Appendix C.
[…]
@@
Number of Hops
Set Operations
Plurality
Moving from expecting a singular answer to allowing multiple valid answers, e.g., “What companies are the [(Org Sub-Role +) Org Role] in the agreement?".
@@ 2.5 Extraction of QA Pairs from Templates
Table 3 illustrates how questions are systematically constructed by incrementally increasing complexity across three dimensions introduced in Section 2.4: number of hops (H), plurality (P), and number of set operations (#SO). Each row in the table corresponds to a specific question template, with associated values for P, H, and #SO, the query path in a KG, and the set operation logic used.
Question instances are constructed by dynamically substituting entities into these templates using SPARQL queries over the knowledge graph described in Section 2.3. This ensures precise and scalable QA pair creation, with answer labels grounded in annotated relationships. While our framework supports arbitrarily complex combinations, we observed that highly compositional templates often become inapplicable to real documents. For instance, there are no individuals in an agreement who hold three distinct positions or no organization with the combination of roles required to fulfill a complex query. Thus, we have to limit template depth since such complex conditions are typically not present in the source documents. The list of all the created templates can be found in Appendix C.
To support evaluation, we define a composite question complexity level L=H+P+#SOL\!=\!H\ \!+\!P\!+\!\#SO, and group QA pairs into three categories. Easy (L=1L\!=\!1): simple, single-hop, singular-answer queries. Medium ($2\!\leqslant\!L\!\leqslant\!4$): questions with moderate complexity via hops, set operations, or plurality. Hard (L=5L\!=\!5): questions with the most complex combinations of all three dimensions.
The resulting dataset, KG-MuLQA-D, comprises 20,139 QA pairs from 170 credit agreements, surpassing most long-context QA benchmarks in size (Table 1). This scale is achieved through efficient question extraction from structured annotations using reusable templates.
@@ 3 Evaluation
In this section, we describe the setup for our experiments. We evaluate multiple long-context LLMs listed in Section 3.1 across different question complexity levels, as defined in Section 2.4. Performance is measured using four metrics detailed in Section D.2. In addition, we conduct human evaluation of LLM responses on a small sample of QA pairs (Section D.3) to discover that F1-score and the LLM-as-a-Judge metrics correlate with human’s expectations most. Our prompting strategy is outlined in Section D.1. The results of our level-based evaluation are presented in Table 4, with the analysis provided in Section 4.